OpenStudy (anonymous):

i will MEDAL PLZ help Solve. 8y – 24 > –16 A. y > –5 B. y > –1 C. y > 1 D. y > 5

OpenStudy (kendricklamar2014):

First Step: Add 24 to both sides:\[8y - 24 + 24 > -16 + 24\] Which equals:\[8y > 8\]

OpenStudy (anonymous):

ok

OpenStudy (kendricklamar2014):

Now, we divide both sides by 8:\[\left(\begin{matrix}8y \\ 8\end{matrix}\right)\left(\begin{matrix}8 \\ 8\end{matrix}\right)\]

OpenStudy (kendricklamar2014):

Which gives you the answer:\[y > 1\]

OpenStudy (anonymous):

(1y) (1)???

OpenStudy (anonymous):

oh ok

OpenStudy (kendricklamar2014):

8y divide by 8 only leaves y because the 8's cancel out

OpenStudy (anonymous):

got it

OpenStudy (kendricklamar2014):

and on the other sides of the equation, 8 divided by 8 leaves you with 1

OpenStudy (anonymous):

ok so what do we do with the 24 and -16

OpenStudy (kendricklamar2014):

The first step we added +24 to both sides which canceled the 24's on the left side of the equation, and -16 + 24 is 8